# When Can LLMs Actually Correct Their Own Mistakes? A Critical Survey of Self-Correction of LLMs

### v1

@@ 5 Self-Correction with External Tools or Knowledge
Although using external tools or knowledge is effective, we raise caution that the way of using external tools or knowledge influences the research questions we can verify (§3.1). As shown in Table 5, some prior studies (Gao et al., 2023; Yu et al., 2023; Zhao et al., 2023) use external knowledge only for self-correction, while they can also directly use external knowledge to improve initial responses. For example, RARR (Gao et al., 2023) uses external knowledge to detect mistakes in initial responses, while it does not use any external knowledge when generating initial responses. These methods are reasonable when only focusing on [RQ3] the performance of final responses, but it is not fair to use them for evaluating [RQ2] whether self-correction can improve from the best-possible initial responses. In contrast, using code interpreters for self-correction (Gou et al., 2024; Chen et al., 2024c) can be regarded as using best-possible initial responses because there is no easy way to improve the initial response generation directly.
@@
| A | B | C | Requirements for Verifying the Target RQs | ||
@@
| ✓ | ✓ | ✓ | When using fine-tuning, reporting the detailed settings, including the number of annotations and computational cost required to achieve the reported performance. | (§6) | Required |
@@
| ✓ | Using external tools or knowledge to improve initial response generation as much as possible. | (§5) | Required | ||
@@
| ✓ | Fine-tuning initial response generators as well, as much as possible. | (§6) | Required | ||
| ✓ | Evaluating the minimum required size of training data that enables self-correction. | (§6) | Recommended | ||
@@
| ✓ | Comparing with strong baselines using comparable computational cost. | (§3.3) | Required |
@@
| When not using external tools or knowledge available in real-world applications, explicitly reporting that the evaluation is done under weak conditions. | (§5) | Required |
| Evaluating with external tools or knowledge available in real-world applications. | (§5) | Recommended |
6 Self-Correction with Fine-tuning
Prior work shows that fine-tuning LLMs for generating feedback or refining responses improves the self-correction capability. A common approach fine-tunes feedback models to generate reference feedback given initial responses and fine-tunes refinement models to generate reference answers given the initial responses and reference feedback (Ye et al., 2023; Lee et al., 2023; Saunders et al., 2022). Frameworks: The first approach fine-tunes the same model to correct its own responses. Most studies fine-tune models for all stages: initial responses, feedback, and refinement (Saunders et al., 2022; Ye et al., 2023; Lee et al., 2023). Another approach corrects responses from larger models using smaller fine-tuned models. This cross-model correction approach often instructs the larger models to refine their own responses using feedback by fine-tuned models (Yang et al., 2022b; Welleck et al., 2023; Akyurek et al., 2023; Paul et al., 2024), which can be viewed as using the small fine-tuned models as external feedback. Training Strategies: A popular approach is supervised fine-tuning, which fine-tunes self-correction modules on human-annotated feedback (Saunders et al., 2022), feedback from stronger models (Ye et al., 2023), or synthetic negative responses (Paul et al., 2024). To avoid the cost of collecting human feedback, self-corrective learning (Welleck et al., 2023) selects model-generated feedback that successfully refines responses as training data and RL4L (Akyurek et al., 2023) uses reinforcement-learning. External Tools: Some works fine-tune models to refine responses given feedback from external tools. Self-Edit (Zhang et al., 2023b) uses the results on test cases evaluated by code executors for code generation, and Baldur (First et al., 2023) uses proof assistants for improving proof generation.
@@ 6 Self-Correction with Fine-tuning
As shown in Table 6, many methods with supervised fine-tuning for feedback generation rely on training data with more than 100K instances. These methods require large-scale human annotations to be implemented on state-of-the-art models. We expect future research to explore approaches that do not require large-scale human annotations (§11).
@@
Some studies (Welleck et al., 2023) apply stronger fine-tuning for self-correction models than initial response generation models, which do not use best-possible initial responses in the available resources (§3.2). These approaches can be used to evaluate [RQ3] the performance of the final responses to compare with other methods but cannot be used to evaluate [RQ2] the improvement from best-possible initial responses.
@@ 7 Summary of Our Analysis
Prior studies widely agree that LLMs can refine their responses given reliable feedback (§5). However, generating reliable feedback on their own responses is still observed to be challenging for LLMs without using additional information (§4), although self-correction is motivated by the hypothesis that recognizing errors is easier than avoiding them (Saunders et al., 2022). We suggest that self-correction research analyze the quality of generated feedback in more detail, not only evaluate refined responses.
@@
Tasks whose verification tasks are much easier than the original tasks (e.g., tasks whose responses are decomposable or verifiable)
Self-Correction with External Information (§5)
@@
Self-Correction with Fine-tuning (§6)
Self-correction works in many tasks when large train data for self-correction is available
@@ 9 Differences from Other Survey
Pan et al. (2024) provide a comprehensive survey on broad studies related to self-correction. Our work specifically focuses on (inference-time) self-correction and provides a more detailed and critical analysis of prior work. Huang et al. (2024) analyze problems in evaluation settings of self-correction research, which motivates our work. They focus on analyzing a few papers on intrinsic self-correction in reasoning tasks. We provide a more comprehensive analysis of self-correction with in-context learning, external tools, and fine-tuning.
@@ 10 Related Work of Self-Correction
of mistakes in LLM responses using LLMs (possibly with external information) has been studied in various domains, including misinformation detection (Zhang et al., 2023c; Chern et al., 2023; Chen and Shu, 2024; Mishra et al., 2024), context-faithfulness (Wang et al., 2020; Durmus et al., 2020; Scialom et al., 2021), harmful content detection (Rauh et al., 2022), and bias detection (Blodgett et al., 2020; Feng et al., 2023).
@@
or self-improvement is an approach to train models using their responses. Some studies use self-evaluation or self-correction for creating training data (Bai et al., 2022; Gulcehre et al., 2023) or use self-evaluation as training signals (Pang et al., 2024). Another approach improves the reasoning of LLMs using LLM-generated reasoning by selecting high-quality outputs using ground-truth final answers (Zelikman et al., 2022) or self-consistency (Huang et al., 2023). As another direction, Meng et al. (2022) use sentences generated by LLMs with high confidence for training classifiers.
@@ 11 Future Directions
Prior studies indicate that it is difficult for LLMs to generate feedback on their own responses with in-context learning (§4, 7). However, most studies in intrinsic self-correction use simple prompts, and there is room for improvement. A possible direction to improve feedback is to apply (reference-free and point-wise) LLM-based evaluation metrics. Recent approaches include using human-written evaluation criteria (Chiang and Lee, 2023; Liu et al., 2023) and decomposing responses (Saha et al., 2023; Min et al., 2023). Another direction is to evaluate the confidence in their responses using generation probabilities (Varshney et al., 2023; Jiang et al., 2023b) or generating new questions from their answers to evaluate logical consistency (Jung et al., 2022; Tafjord et al., 2022; Wu et al., 2024).
@@
Difficulty of self-evaluation differs from task to task (§4), while many studies assume that verification is consistently easier than generation (e.g., Saunders et al., 2022). We expect that there are unexplored tasks in which intrinsic self-correction works well. For example, LLM-based evaluation is often studied in open-ended text generation, such as dialogue generation (Fu et al., 2023; Liu et al., 2023), suggesting that reasonable model-based feedback is available.
Fine-tuning with Small Training Data.
Fine-tuning of feedback generation often relies on large training data, which requires large-scale human annotations (§6). We expect future work to explore self-correction with smaller training data. Although reinforcement learning (Akyurek et al., 2023) or self-corrective learning (Welleck et al., 2023) do not require human feedback, they require reasonable reward functions for evaluating LLM responses, which are not available in many tasks. For example, RL4F (Akyurek et al., 2023) uses ROUGE as a reward function for text summarization and action planning, which is sub-optimal.
@@
Prior studies show that large-scale fine-tuning on reference feedback improves the self-correction capability of LLMs (§6). This observation suggests that the current approach or datasets for pre-training LLMs are insufficient to make LLMs acquire self-correction capability. We expect future work to explore pre-training strategies to improve the intrinsic self-correction capability of LLMs.
@@ Acknowledgments
This work was supported by a Cisco Research Grant.

### v2

@@ 5.1 Self-Correction with External Tools or Knowledge
Although using external tools or knowledge is known to be effective in self-correction, we raise caution that the way of using external tools or knowledge influences the research questions we can verify (§3.1). As shown in Table 5, some prior studies (Gao et al., 2023; Yu et al., 2023; Zhao et al., 2023) use external knowledge only for self-correction, while they can also directly use external knowledge to improve the initial response generation process. For example, RARR (Gao et al., 2023) uses external knowledge to detect mistakes in initial responses, while it does not use any external knowledge when generating initial responses. These methods are reasonable when only focusing on [RQ3] the performance of final responses, but it is not fair to use them for evaluating [RQ2] whether self-correction can improve from the best-possible initial responses. In contrast, using code interpreters for self-correction (Gou et al., 2024; Chen et al., 2024d) can be regarded as using best-possible initial responses because there is no easy way to improve the initial response generation directly.
Verifiable Tasks.
Some tasks have a property that allows the correctness of the responses to be verified easily, even without external information. For example, the constrained generation task evaluated in Self-Refine (Madaan et al., 2023) is a task to generate a sentence that includes five specified words. We can easily evaluate the correctness by checking whether the five words are included in the generated sentence. Tree-of-thought (Yao et al., 2023) is a generate-and-rank method for verifiable tasks,111Tree-of-thought is a generate-and-rank method and not a self-correction method in our definition. such as Game of 24, the task to obtain 24 using basic arithmetic operations () and provided four integers. For Game of 24, we can easily verify the answer by checking whether the generated answer is 24. We consider self-correction to work well in these tasks because they are in the same situations as using strong external tools or the oracle information to generate feedback.
@@
| RQ1 | RQ2 | RQ3 | Requirements for Verifying the Target RQs | ||
@@
| ✓ | ✓ | ✓ | When using fine-tuning, reporting the detailed settings, including the number of annotations and computational cost required to achieve the reported performance. | (§5.2) | Required |
@@
| ✓ | Using external tools or knowledge to improve initial response generation as much as possible. | (§5.1) | Required | ||
@@
| ✓ | Fine-tuning initial response generators as well, as much as possible. | (§5.2) | Required | ||
| ✓ | Evaluating the minimum required size of training data that enables self-correction. | (§5.2) | Recommended | ||
@@
| ✓ | Comparing with strong baselines using comparable computational cost. | (§6) | Required |
@@
| When not using external tools or knowledge available in real-world applications, explicitly reporting that the evaluation is done under weak conditions. | (§5.1) | Required |
| Evaluating with external tools or knowledge available in real-world applications. | (§5.1) | Recommended |
5.2 Self-Correction with Fine-tuning
Prior work shows that fine-tuning LLMs for generating feedback or refining responses improves the self-correction capability. A common approach fine-tunes feedback models to generate reference feedback given initial responses and fine-tunes refinement models to generate reference answers given the initial responses and reference feedback (Ye et al., 2023; Lee et al., 2024; Saunders et al., 2022). Frameworks: The first approach fine-tunes the same model to correct its own responses. In this approach, most methods fine-tune models for all stages: initial responses, feedback, and refinement (Saunders et al., 2022; Ye et al., 2023; Lee et al., 2024). Another approach corrects responses from larger models using smaller fine-tuned models. This cross-model correction approach often instructs the larger models to refine their own responses using feedback from the smaller fine-tuned models (Yang et al., 2022b; Welleck et al., 2023; Akyurek et al., 2023; Paul et al., 2024), which can be viewed as using the small fine-tuned models as external tools. Training Strategies: A popular approach is supervised fine-tuning, which fine-tunes self-correction modules on human-annotated feedback (Saunders et al., 2022), feedback from stronger models (Ye et al., 2023), or synthetic negative responses (Paul et al., 2024). As other approaches, to avoid the cost of collecting human feedback, self-corrective learning (Welleck et al., 2023) selects model-generated feedback that successfully refines responses as training data, and RL4L (Akyurek et al., 2023) uses reinforcement-learning. External Tools: Some works fine-tune models to refine responses given feedback from external tools. Self-Edit (Zhang et al., 2023b) uses the results on test cases evaluated by code executors for code generation, and Baldur (First et al., 2023) uses proof assistants for improving proof generation.
@@ 5.2 Self-Correction with Fine-tuning
As shown in Table 6, many methods with supervised fine-tuning for feedback generation rely on training data with more than 100K instances. These studies often use feedback generated by stronger models to simulate human annotation, but this approach requires large-scale human annotations to be implemented on state-of-the-art models. We expect future research to explore approaches that do not require large-scale human annotations (§11).
@@
Some studies (Welleck et al., 2023) apply stronger fine-tuning for self-correction models than initial response generation models, which do not use best-possible initial responses in the available resources (§3.2). This approach can be used to evaluate [RQ3] the performance of the final responses to compare with other methods but cannot be used to evaluate [RQ2] the improvement from best-possible initial responses.
6 Strong Baselines
Self-correction involves multiple LLM calls for generating feedback and refinement. Therefore, to claim that [RQ3] the performance of the final outputs from self-correction frameworks is better than other approaches, it should be compared with sufficiently strong baselines, possibly relying on additional LLM calls or computational cost. Many self-correction studies do not compare their methods with strong baselines, although some studies pointed out this issue and compare self-correction with self-consistency (Gou et al., 2024; Huang et al., 2024a) or pass@k in code generation (Zhang et al., 2023b; Olausson et al., 2024). We encourage future research to compare self-correction with strong baselines, including self-consistency and generate-and-rank, to further explore RQ3.
Self-Consistency
(Wang et al., 2023) is an approach that generates multiple responses for the same input and takes the majority vote of the final answers in reasoning tasks. The idea of selecting good responses using the consistency between multiple responses from the same model has also been extended to other tasks such as text generation (Manakul et al., 2023; Elaraby et al., 2023; Chen et al., 2024c) and code generation (Shi et al., 2022).
Generate-and-Rank
is an approach that generates multiple responses and selects the best response using verifiers. Post-hoc approach ranks responses using self-evaluation (Weng et al., 2023; Zhang et al., 2023d), confidence (Manakul et al., 2023), fine-tuned verifiers (Cobbe et al., 2021; Shen et al., 2021; Lightman et al., 2024), or verifiers with external tools (Shi et al., 2022; Chen et al., 2023a; Ni et al., 2023). Feedback-guided decoding generates multiple responses and selects the best response for each reasoning step using generation probability (Hao et al., 2023; Tyen et al., 2024), prompted self-evaluation (Jung et al., 2022; Creswell and Shanahan, 2022; Xie et al., 2023; Yao et al., 2023; Miao et al., 2024), or fine-tuned verifiers (Uesato et al., 2022; Tafjord et al., 2022; Yang et al., 2022a; Asai et al., 2024).
@@ 7 Summary of Our Analysis
Prior studies widely agree that LLMs can refine their responses given reliable feedback (§5). However, generating reliable feedback on their own responses is still observed to be challenging for LLMs without using additional information (§4). In other words, for the current LLMs, the hypothesis that recognizing errors is easier than avoiding them (Saunders et al., 2022) is only true for certain tasks whose verification is exceptionally easy, according to our analysis of the experiments in prior studies. We recommend that self-correction research analyze the quality of generated feedback in more detail, not only evaluate the downstream performance of the refined responses.
@@
Tasks whose verification tasks are much easier than the original tasks (e.g., tasks whose responses are decomposable)
Self-Correction with External Information (§5.1)
@@
Self-Correction with Fine-tuning (§5.2)
Self-correction works in many tasks when large training data for feedback generation is available
@@ 9 Differences from Other Survey
Pan et al. (2024) provide a comprehensive survey on broad topics related to self-correction, including training strategies. Our work specifically focuses on (inference-time) self-correction and provides a more detailed and critical analysis of prior work. Huang et al. (2024a) provide an analysis of problems in the evaluation settings of self-correction research, which motivates our work. They focus on analyzing a few papers on intrinsic self-correction in reasoning tasks. We provide a more comprehensive analysis of self-correction with in-context learning, external tools, and fine-tuning.
@@ 10 Related Work of Self-Correction
of mistakes in LLM responses using LLMs (possibly with external information) has been studied in various domains, including misinformation detection (Zhang et al., 2023c; Chern et al., 2023; Chen and Shu, 2024; Mishra et al., 2024), context-faithfulness (Wang et al., 2020; Durmus et al., 2020; Scialom et al., 2021), harmful content detection (Rauh et al., 2022), and bias detection (Blodgett et al., 2020; Feng et al., 2023). However, recent studies (Tyen et al., 2024; Kamoi et al., 2024) show that even strong LLMs often cannot detect their own mistakes in various tasks.
@@
or self-improvement is an approach to train models using their own responses. Some studies use self-evaluation or self-correction for creating training data (Bai et al., 2022; Gulcehre et al., 2023) or use self-evaluation as training signals (Pang et al., 2024). Another approach improves the reasoning of LLMs using LLM-generated reasoning by selecting high-quality outputs using ground-truth final answers (Zelikman et al., 2022) or self-consistency (Huang et al., 2023). As another direction, Meng et al. (2022) use sentences generated by LLMs with high confidence for training classifiers.
@@ 11 Future Directions
Prior studies indicate that it is difficult for LLMs to generate feedback on their own responses with in-context learning (§4, 7). However, most studies in intrinsic self-correction (Madaan et al., 2023; Huang et al., 2024a) use simple prompts for generating feedback, and there is room for improvement. A possible direction to improve feedback is to apply (reference-free and point-wise) LLM-based evaluation metrics. Recent approaches for improving the model-based evaluation include using human-written evaluation criteria (Chiang and Lee, 2023; Liu et al., 2023) and decomposing responses (Saha et al., 2024; Min et al., 2023). As another direction, recent studies in self-correction propose frameworks using the confidence in their responses, estimated by generation probabilities (Varshney et al., 2023; Jiang et al., 2023b), prompting (Li et al., 2024a), or generating new questions from their answers to evaluate logical consistency (Jung et al., 2022; Tafjord et al., 2022; Wu et al., 2024).
@@
The difficulty of self-evaluation differs from task to task (§4), while many studies assume that verification is consistently easier than generation. We expect that there are unexplored tasks in which intrinsic self-correction works well, although self-correction research mostly focuses on reasoning tasks such as math reasoning and coding (Madaan et al., 2023; Gou et al., 2024; Huang et al., 2024a). For example, LLM-based evaluation is often studied in open-ended text generation, such as dialogue generation and text summarization (Fu et al., 2024; Liu et al., 2023), suggesting that reasonable model-based feedback is available for these tasks.
Fine-tuning on Small Training Data.
Fine-tuning of feedback generation often relies on large training data, which requires large-scale human annotations (§5.2). We expect future work to explore self-correction with smaller training data. Although reinforcement learning (Akyurek et al., 2023) or self-corrective learning (Welleck et al., 2023) do not require human feedback, they require reasonable reward functions for evaluating LLM responses, which are not available in many tasks. For example, RL4F (Akyurek et al., 2023) uses ROUGE as a reward function for text summarization and action planning, which is sub-optimal.
@@
Prior studies show that large-scale fine-tuning on reference feedback improves the self-correction capability of LLMs (§5.2). This observation suggests that the current approach or datasets for pre-training LLMs are insufficient to make LLMs acquire self-correction capability. We expect future work to explore pre-training strategies to improve the intrinsic self-correction capability of LLMs.
12 Conclusion
We provide a critical survey of self-correction to identify in which conditions LLMs can self-correct their mistakes. Our analysis reveals that many studies fail to define their research questions clearly or design experiments appropriately. To tackle these issues, we categorize research questions and frameworks in self-correction research and provide a checklist for conducting appropriate experiments.
@@ Acknowledgments
This work was supported by a Cisco Research Grant. We appreciate valuable suggestions from the action editor and anonymous reviewers.